# Hardware-efficient Customized Keyword Spotting with Spectral-Temporal Graph Attentive Pooling and a hybrid Loss

### v1

@@ Abstract
Most of the existing systems designed for keyword spotting (KWS) rely on a predefined set of keyword phrases. However, the ability to recognize customized keywords is crucial for tailoring interactions with intelligent devices. In this paper, we present a novel framework for customized KWS. This framework leverages the hardware-efficient LiCoNet architecture as the encoder, enhanced by a spectral-temporal pooling layer and a hybrid loss function to facilitate effective word embedding learning. The experimental results on a substantial internal dataset have demonstrated the distinct advantages of the proposed framework. LiCoNet performs at a similar level ($1.98\%$ FRR at $0.3$ FAs/Hr) to the computationally intensive Conformer, which requires 13x computational resources.
Index Terms—  Query-by-example Keyword Spotting, Conformer, LicoNet, Spectral-temporal Attentive Pooling, AAM, SoftTriplet
@@ 1 Introduction
A keyword spotting (KWS) system serves the purpose of detecting a predetermined keyword within a continuous real-time audio stream. This capability is pivotal in facilitating interactions between users and voice assistants. The introduction of a customized KWS system, which empowers users to define their own keywords, offers a substantial degree of flexibility and personalization in user experiences. However, this customization also presents significant challenges, such as the need for a small KWS memory footprint, minimizing latency, and handling user-defined keyword phrases that may not align with the training data distribution.
One approach to address these challenges involves the use of Query-by-Example (QbyE) techniques [1]. In this context, the KWS system utilizes audio samples of keywords provided by users to generate fixed-length embeddings. These embeddings are then employed to assess the similarity between test samples and the enrolled keywords within the embedding space, ultimately determining the presence of a keyword. The uppermost diagram in Figure 1 illustrates a broad framework for QbyE KWS. Within this framework, there is a pooling layer positioned after the encoder, which is responsible for creating an information-rich embedding. This embedding is subsequently supplied to a classifier to distinguish between sub-words or words.
In previous studies, transformers have found extensive use in encoder modeling because of their substantial modeling capabilities [2] [3]. Nevertheless, attention-based models are associated with significant computational demands and impose a high runtime memory burden when deployed on hardware. This characteristic makes them unsuitable for an always-on KWS system. The linearized convolution network (LiCoNet) as introduced in [4] for KWS modeling, offers excellent hardware efficiency while maintaining a high level of model effectiveness.
@@
In this study, we introduce a LiCoNet-based, hardware-efficient framework for customized KWS modeling. We employ spectro-temporal graph attentive pooling (GAP) [5] to generate informative embeddings. This pooling layer demonstrates strong capability in comprehending the complex relationships within spectral-temporal data. During the training phase, we formulate a hybrid loss function that combines elements from the Additive Angular Margin (AAM) and SoftTriplet losses, which are widely employed in tasks such as face recognition [6] and speaker recognition [7]. The hybrid loss is crafted to enhance the distinctiveness of words and phonemes while simultaneously reducing the variability in learned embeddings attributed to speakers. Our experimental results, conducted on a substantial internal dataset, showcase the advantages of our proposed framework, which features GAP and the hybrid loss, for customized KWS. Notably, LiCoNet achieves performance levels similar to those of the computationally intensive Conformer, which requires $13$x computational resources.
@@ 2.1 Encoder-decoder Architecture
Fig. 1: The customized KWS training framework.
@@ 2.2.1 ECAPA_TDNN
The entire training and testing process shares similarities with the speaker verification (SV) task. Consequently, we consider ECAPA_TDNN, a commonly used backbone model architecture for the SV task [8], as a potential choice for the acoustic feature encoder in this study. The ECAPA_TDNN model consists of a 1D convolution followed by three 1D SE-Res2Blocks, 1D convolution, attentive statistical pooling, and a fully connected (FC) layer. After each layer within the SE-Res2Block, we apply non-linear ReLU activation and batch normalization (BN). The embedding feature vectors are extracted from the FC layer.
2.2.2 Convolution-augmented Transformer (Conformer)
The Conformer architecture has proven its remarkable effectiveness within the sequence-to-sequence domain [9] and has achieved significant success in the realm of speech recognition tasks [10] [11] [12]. This architecture seamlessly integrates the capabilities of both convolutional and self-attention mechanisms, providing a flexible and exceptionally potent solution for learning feature representations from sequential data. Each Conformer block comprises four consecutive modules, including a feed-forward module, a self-attention module, a convolution module, and a second feed-forward module [9]. This Conformer-based encoder demonstrates the ability to leverage position-specific local features, facilitated by the convolution module, while simultaneously capturing content-based global interactions through the self-attention module.
2.2.3 Linearized Convolution Network (LicoNet)
LiCoNet represents a hardware-efficient architecture specifically designed for the KWS task, as detailed in [4]. This architecture is carefully crafted as a streaming convolution network, using equivalent linear operators to ensure efficient inference while preserving a high level of detection accuracy. Each LiCo-Block is structured as a bottleneck configuration composed of three 1D convolution layers. The initial layer employs streaming convolution with a kernel size greater than $1$, followed by two subsequent point-wise convolutions.
@@ 2.3 Feature Aggregator
Pooling plays an essential role in neural architectures, serving the purpose of distilling crucial insights from sequential data while preserving essential contextual details. In the context of our study, we investigate two distinct pooling strategies for word embedding learning.
Attentive Statistic Pooling (ASP) ASP combines the strengths of both statistical pooling and attention mechanisms [13]. Attention allows the model to dynamically weigh the importance of different elements along the temporal dimension, enabling the extraction of salient features that are crucial for the task at hand.
Spectral-temporal Graph Attentive Pooling (GAP) GAP has gained success in the field of speech and audio processing [5] [14]. It leverages the power of graph neural networks to comprehend complex relationships within spectral-temporal data. The spectral and temporal attention module comprises three graph attention blocks, each housing the graph attention network (GAT) and graph pooling. This configuration empowers the model to adapt the pooling procedure dynamically, facilitating the extraction of crucial features.
@@ 2.4 Loss Function
2.4.1 Additive Angular Margin (AAM)
The AAM loss is designed to enhance the discrimination power of neural networks by emphasizing the angular separation between class embeddings and is prevalent in the context of face recognition and feature embedding [6]. The loss function is defined as,
$\small\mathcal{L}_{aam}=-\mathop{log}\frac{e^{\textbf{\em s}\>cos(\theta_{y_{i}}+m)}}{e^{\textbf{\em s}\>cos(\theta_{y_{i}}+m)}+\sum_{j=1,j\neq y_{i}}^{C}e^{\textbf{\em s}\>cos(\theta_{j})}},$ (1)
where $\theta_{j}$ is the angle between the feature $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and the weight $\mathbf{w}_{j}\in\mathbb{R}^{d}$. $\mathbf{w}_{j}$ denotes the $j$-th column of the weight $[\mathbf{w}_{1},\cdots,\mathbf{w}_{C}]\in\mathbb{R}^{d\times C}$ of the last fully-connected layer that maps $d$-dimensional embeddings to the logits. $C$ is the number of classes. s is a rescaling factor. An additive angular margin $m$ is applied for adjustment.
@@ 2.4.1 Additive Angular Margin (AAM)
Fig. 2: DET curves of Conformer using various loss formulations (left) and of different encoders using the hybrid loss (right).
2.4.2 SoftTriplet
For customized KWS tasks, it’s important to note that the training and testing datasets have no overlap in data distribution. This uniqueness necessitates that the model possesses strong generalization capabilities. The QbyE system can hence be conceptualized as an optimization problem featuring triplet constraints. The primary objective is to minimize the distance between the embedding of an enrolled keyword and those of the same word phrase while simultaneously maximizing the distance between embeddings of different word phrases. To address this optimization goal and simultaneously account for intra-class variance—such as speaker variability or variations in speaking rates—we have chosen to adopt the SoftTriplet loss [15] as a word-level loss function,
$\small\mathcal{S}^{\prime}_{i,c}=\sum_{k}\frac{exp(\frac{1}{\gamma}\mathbf{x}_{i}^{\top}\mathbf{w}_{c}^{k})}{\sum_{k}exp(\frac{1}{\gamma}\mathbf{x}_{i}^{\top}\mathbf{w}_{c}^{k})}\mathbf{x}_{i}^{\top}\mathbf{w}_{c}^{k}$ (2)
@@ 2.4.2 SoftTriplet
$\small\mathcal{L}_{st}({\mathbf{x}}_{i})=-\mathop{log}\frac{exp(\lambda(\mathcal{S}^{\prime}_{i,y_{i}}-\delta))}{exp(\lambda(\mathcal{S}^{\prime}_{i,y_{i}}-\delta))+\sum_{j}exp(\lambda\mathcal{S}^{\prime}_{i,j})},$ (3)
@@
where $\mathcal{S}^{\prime}_{i,c}$ is the similarity between feature $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and the class $c$. $\mathbf{w}_{c}^{k}$ is the $k$-th center out of $K$ centers for the class $c$. $\delta$ is a predefined margin. $\lambda$ denotes a scaling factor.
2.4.3 Hybrid Loss
Phoneme Context Phonemes serve as the fundamental phonetic units that compose spoken words. Incorporating the context of phonemes into modeling offers a nuanced source of information for refining word embeddings. In our approach, as depicted in Fig. 1, we introduce a dedicated phoneme classifier into the training framework. The phoneme loss is computed by aggregating the frame-level AAM loss, applied to phoneme labels, across all frames.
@@ 2.4.3 Hybrid Loss
Speaker Variability Acoustic variations related to individual speakers, such as differences in pitch, tone, or pronunciation, exert a significant influence on speech modeling. Existing approaches in QbyE KWS often assume that the system user is the same as the enrolled speaker. To address and disentangle speaker dependency within the application, we have incorporated a reverse speaker loss into our methodology, with the objective of learning speaker-independent embeddings (see Fig. 1). More specifically, we have devised an AAM-based reverse speaker loss, which is employed to maximize the speaker classification loss through the application of a gradient reversal layer (GRL) [16] during the training process. The parameter-free GRL functions as an identity transform during forward propagation but reverses gradients during back-propagation, feeding them into the preceding layer.
@@
Consequently, our hybrid loss function is constructed as a combination of word-level loss, phoneme-level loss, and the reverse speaker loss.
$\small\begin{split}\mathcal{L}(\mathbf{x},\mathbf{y})=\mathcal{L}_{st}\left(\mathbf{x},y^{w}\right)-\eta\mathcal{L}_{aam}(\mathbf{x},y^{s})+\mu\mathcal{L}_{aam}(\mathbf{x},y^{p}),\end{split}\vspace{-2ex}$ (4)
@@
where $\mathbf{x}$ is the acoustic feature vector, $\mathbf{y}=(y^{w},y^{s},y^{p})$ $y^{w}$ is the word label, $y^{s}$ is the speaker label, $y^{p}\in\mathbb{R}^{T}$ is the phoneme label sequence, $\eta$ and $\mu$ are scaling factors.
@@ 3.1 Dataset
We use the Librispeech [17] dataset containing 960 hours of read English audiobooks sampled at $16$ kHz along with transcriptions. We employ a pre-trained acoustic model for the force-alignment to segment utterances into individual words. Each word-level segment is standardized to $2$s long by clipping or zero padding on both sides of the audio.
We use the internal aggregated and de-identified keyword dataset for evaluation. The positive data contains $275.7$k utterances from $629$ speakers. The total duration of negative data is up to $200$ hours. We extract acoustic features using $40$-dimensional log Mel-filterbank energies computed over a $25$ms window every $10$ms.
@@ 3.2 Experimental Setup
Model architecture We conduct the experiments on three model types: ECAPA_TDNN, Conformer, and LicoNet. We setup ECAPA_TDNN with $128$ channels in the convolution layers and a $64$ dimensional bottleneck in the SE-Block and attention module. The scale dimension $s$ in the Res2Block is $8$. Conformer has two heads per multi-headed self-attention layer with $128$ input and output nodes [18]. The linear hidden units have a dimensionality of $192$, and the convolution module uses the kernel size of $7$. We construct LiCoNet by stacking $5$ LiCo-Blocks with the expansion factor of $6$ and the kernel size of $5$ [4]. Table 2 presents the model size and floating point operations per second (FLOPs) of $2$s audio for each encoder model.
@@
Feature aggregator We compare Graph Attentive Pooling (GAP) against Attentive Statistic Pooling (ASP) as the feature aggregator. The spectral, temporal and spectro-temporal attention blocks use pooling ratios of $0.71$, $0.86$, and $0.71$, respectively.
Loss function We focus on investigating the effectiveness of Additive Angular Margin (AAM) and SoftTriplet in customized KWS modeling, with cross-entropy as the baseline. The AAM Softmax margin $m$ and scale s in Eq. 1 are set to $0.2$ and $32$, respectively. The SoftTriplet loss uses the scaling factor $\lambda=60$, $\gamma=1$, the margin $\delta=0.03$, and $K=10$. $\eta$ and $\mu$ in Eq. 4 is set to $0.1$ and $0.5$, respectively.
@@
Training and testing protocols All KWS models are trained to predict 1002 targets (i.e., the top $1$k frequent words, Silence, and unknown words). We use a batch size of 64 with 8 GPUs for 40-epoch training. We adopt the triangular2 policy as described in [19] in conjunction with the Adam optimizer with a cyclical learning rate increased from $1$e-$8$ to $1$e-$3$ in $20$k warming-up updates. During testing, $3$ utterances were randomly picked as enrollments. For a given query, the cosine distance is used to compare the similarity between the query embedding and the 3 enrolled ones. The minimum distance is used to compare against a threshold value to make the detection decision. We present the model performance by plotting detection error trade-off (DET) curves, where the x-axis and y-axis represent the number of false accepts (FA) per hour and false reject rate (FRR), respectively.
4 Results and Discussion
Model Performance

### v2

# Query-by-Example Keyword Spotting Using Spectral-Temporal Graph Attentive Pooling and Multi-Task Learning

@@ Abstract
Existing keyword spotting (KWS) systems primarily rely on predefined keyword phrases. However, the ability to recognize customized keywords is crucial for tailoring interactions with intelligent devices. In this paper, we present a novel Query-by-Example (QbyE) KWS system that employs spectral-temporal graph attentive pooling and multi-task learning. This framework aims to effectively learn speaker-invariant and linguistic-informative embeddings for QbyE KWS tasks. Within this framework, we investigate three distinct network architectures for encoder modeling: LiCoNet, Conformer and ECAPA_TDNN. The experimental results on a substantial internal dataset of $629$ speakers have demonstrated the effectiveness of the proposed QbyE framework in maximizing the potential of simpler models such as LiCoNet. Particularly, LiCoNet, which is 13x more efficient, achieves comparable performance to the computationally intensive Conformer model ($1.98\%$ vs. $1.63\%$ FRR at $0.3$ FAs/Hr).
keywords:
Query-by-example Keyword Spotting, Conformer, LicoNet, Spectral-temporal Attentive Pooling, Additive Angular Margin, SoftTriplet
@@ 1 Introduction
A keyword spotting (KWS) system serves the purpose of detecting a predetermined keyword within a continuous real-time audio stream. This capability is pivotal in facilitating interactions between users and voice assistants. The introduction of a customized KWS system, which empowers users to define their own keywords, offers a substantial degree of flexibility and personalization in user experiences. However, in addition to the typical difficulties associated with KWS, such as maintaining a small memory footprint and minimizing latency, the customization process also introduces a significant challenge: user-defined keyword phrases may not align with the distribution of the training data, resulting in inferior detection performance. One approach for customized KWS involves the use of Query-by-Example (QbyE) techniques [1]. In this context, the KWS system utilizes audio samples of keywords provided by users to generate fixed-length embeddings. These embeddings are then employed to assess the similarity between test samples and enrolled keywords within the embedding space, ultimately determining the presence of a keyword.
Extensive research has explored the application of various neural network architectures in the context of pre-defined KWS tasks. In [2] [3], an encoder-decoder model structure has been proposed, where an acoustic encoder generates senone-level posteriors, and a decoder is independently trained to interpret the encoder outputs as keyword phrases. Following the trend of end-to-end modeling, researchers have also suggested building an end-to-end KWS system that directly outputs detection without distinguishing the encoder and decoder in the training [4] [5]. Despite achieving high detection performance, training a dedicated model for predefined keywords typically requires substantial target data to ensure effectiveness.
Customized KWS systems alleviate data requirements and offer flexibility by supporting keywords beyond a predefined set. Early endeavors on QbyE KWS tasks rely on a pre-trained Automated Speech Recognition (ASR) system. Specifically, the acoustic model of an ASR first generates phonetic posteriors for an audio stream, and dynamic time warping (DTW) is employed to measure the similarity between the posterior sequences of enrolled keywords and testing samples [6] [7] [8]. The CTC-based KWS approach extends this framework by employing CTC forwarding to evaluate the likelihood of each keyword hypothesis given an audio sample. It then generates a final detection score by aggregating the scores from the N-best list [9] [10].
Chen et. al have proposed LSTM-based encoder modeling for learning acoustic embeddings of customized keywords and have shown promising results [1]. [11] further improves encoder modeling by incorporating multi-head attention for feature extraction, using a normalized multi-head attention module for feature aggregation, and integrating SoftTriplet loss to enhance discrimination capabilities, resulting in improved performance and capabilities. However, attention-based models are accompanied by significant computational demands, posing a high runtime memory burden when deployed on hardware. This characteristic makes them unsuitable for an always-on KWS system. To address this challenge, Huang et. al have made an initial attempt by replacing the attention mechanism with the MLPMixer architecture in the QbyE KWS tasks [12]. Despite the improved performance and efficiency compared to ViT, the frequent utilization of matrix transpose operations can still cause significant computation overhead on hardware.
@@
In this study, we introduce an effective framework for QbyE KWS. The framework employs spectro-temporal graph attentive pooling (GAP) [13] and multi-task learning to facilitate informative embedding learning. GAP demonstrates strong capability in comprehending complex relationships within spectral-temporal data. The multi-task learning is designed to minimize the loss of classifying words and phonemes but to maximize the loss of distinguishing speakers, aiming to enhance the distinctiveness of words and phonemes while simultaneously reducing the speaker variability in learned embeddings. The loss function incorporates Additive Angular Margin (AAM) loss and SoftTriplet loss, both widely employed in tasks such as speaker recognition [14] and face recognition [15]. Within this framework, we investigate three distinct network architectures for encoder modeling: LiCoNet, Conformer and ECAPA_TDNN. Our experimental results, conducted on a substantial internal dataset of $629$ speakers, showcase the effectiveness of the proposed QbyE framework in maximizing the potential of simpler models. Notably, within this framework, LiCoNet, which is $13$x more efficient, achieves comparable performance to the computationally intensive Conformer model.
@@ 2.1 Encoder-Decoder Architecture
Figure 1: The customized KWS training framework.
@@ 2.2 Feature Encoder
The primary objective of the encoder is to efficiently learn effective embeddings for both acoustic and linguistic information. In this section, we investigate three distinct network architectures for encoder modeling: LiCoNet, which strikes a balance between model effectiveness and efficiency; Conformer, renowned for its exceptional capabilities in sequence modeling; and ECAPA_TDNN, a widely used backbone for speaker verification (SV).
2.2.1 Linearized Convolution Network (LiCoNet)
LiCoNet represents a hardware-efficient architecture specifically designed for the KWS tasks, as detailed in [16]. This architecture is carefully crafted as a streaming convolution network built upon the MobileNetV2 backbone [17] for training. It uses equivalent linear operators to ensure efficient inference while preserving a high level of detection accuracy. Each LiCo-Block is structured as a bottleneck configuration composed of three 1D convolution layers. The initial layer employs streaming convolution with a kernel size greater than $1$, followed by two subsequent point-wise convolutions.
2.2.2 Convolution-augmented Transformer (Conformer)
The Conformer architecture has proven its remarkable effectiveness within the sequence-to-sequence domain [18] and has achieved significant success in the realm of speech recognition tasks [19] [20] [21]. This architecture seamlessly integrates the capabilities of both convolutional and self-attention mechanisms, providing a flexible and exceptionally potent solution for learning feature representations from sequential data. Each Conformer block comprises four consecutive modules, including a feed-forward module, a self-attention module, a convolution module, and a second feed-forward module [18]. This Conformer-based encoder demonstrates the ability to leverage position-specific local features through the convolution module, while simultaneously capturing content-based global interactions through the self-attention module.
@@ 2.2.3 ECAPA_TDNN
The QbyE KWS training and testing process shares similarities with SV. We hence consider ECAPA_TDNN, a commonly used backbone model architecture for the SV tasks [22], as a potential choice for encoder modeling in this study. The ECAPA_TDNN model consists of a 1D convolution followed by three 1D SE-Res2Blocks, a 1D convolution, an attentive statistical pooling, and a fully connected (FC) layer. After each layer within the SE-Res2Block, we apply non-linear ReLU activation and batch normalization (BN). The embedding feature vectors are extracted from the FC layer.
@@ 2.3 Feature Aggregator
Pooling plays an essential role in neural architectures (see Figure 1), serving the purpose of distilling crucial insights from sequential data while preserving essential contextual details. In the context of our study, we investigate two distinct pooling strategies for word embedding learning.
Attentive Statistic Pooling (ASP) combines the strengths of both statistical pooling and attention mechanisms [23]. Attention allows the model to dynamically weigh the importance of different elements along the temporal dimension, enabling the extraction of salient features that are crucial for the task at hand.
Spectral-temporal Graph Attentive Pooling (GAP) has gained success in the field of speech and audio processing [13] [24]. It leverages the power of graph neural networks to comprehend complex relationships within spectral-temporal data. The spectral and temporal attention module comprises three graph attention blocks, each housing the graph attention network and graph pooling. This configuration empowers the model to adapt the pooling procedure dynamically, facilitating the extraction of crucial features.
@@ 2.4 Loss Function
The learning process of the encoder is designed to maximize the discriminative power of audio embeddings across distinct words [11] [12]. In this study, we propose multi-task learning that not only considers word-level discrimination but also incorporates fine-grained phoneme information and speaker variability, enhancing the overall modeling effectiveness (see Figure 1).
@@
Figure 2: DET curves of Conformer using various loss formulations (left) and of different encoders using the hybrid loss (right).
2.4.1 Word Discrimination
For QbyE KWS tasks, it’s important that the model possesses strong generalization capabilities. The QbyE system can hence be conceptualized as an optimization problem focusing on minimizing the distance between the embeddings of the same word while simultaneously maximizing the embedding distance of different words. We explore two popular losses for this purpose: additive angular margin (AAM) and SoftTriplet.
AAM emphasizes the angular separation between class embeddings and is prevalent in the context of face recognition and feature embedding [15]. The loss function is defined as,
$\small\mathcal{L}_{aam}=-\mathop{\log}\frac{e^{\textbf{\em s}\>\cos(\theta_{y_{i}}+m)}}{e^{\textbf{\em s}\>\cos(\theta_{y_{i}}+m)}+\sum_{j=1,j\neq y_{i}}^{C}e^{\textbf{\em s}\>\cos(\theta_{j})}},$ (1)
where $\theta_{j}$ is the angle between the feature $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and the weight $\mathbf{w}_{j}\in\mathbb{R}^{d}$. $\mathbf{w}_{j}$ denotes the $j$-th column of the weight $[\mathbf{w}_{1},\cdots,\mathbf{w}_{C}]\in\mathbb{R}^{d\times C}$ of the last fully-connected layer that maps $d$-dimensional embeddings to the logits. $C$ is the number of classes and s is a rescaling factor. An additive angular margin $m$ is applied for adjustment.
SoftTriplet has showcased effectiveness for QbyE KWS modeling in [11]. It combines triplet loss and softmax loss, defined as [25],
@@ 2.4.1 Word Discrimination
$\small\mathcal{L}_{st}({\mathbf{x}}_{i})=-\mathop{\log}\frac{\exp(\lambda(\mathcal{S}^{\prime}_{i,y_{i}}-\delta))}{\exp(\lambda(\mathcal{S}^{\prime}_{i,y_{i}}-\delta))+\sum_{j}\exp(\lambda\mathcal{S}^{\prime}_{i,j})},$ (2)
@@
where $\mathcal{S}^{\prime}_{i,c}$ is the similarity between feature $\mathbf{x}_{i}\in\mathbb{R}^{d}$ and the class $c$. $\delta$ is a predefined margin. $\lambda$ denotes a scaling factor.
2.4.2 Speaker Variability
@@
Acoustic variations related to individual speakers, such as differences in pitch, tone, or pronunciation, exert a significant influence on speech modeling. Existing approaches in QbyE KWS often assume that the system user is the same as the enrolled speaker. To disentangle speaker dependency within the application, we incorporate a reverse speaker loss into our methodology, with the objective of learning speaker-independent embeddings. Specifically, we design an AAM-based reverse speaker loss, which is employed to maximize the speaker classification loss through the application of a gradient reversal layer (GRL) [26] during the training process. The parameter-free GRL functions as an identity transform during forward propagation but reverses gradients during back-propagation.
2.4.3 Phoneme Context
Phonemes serve as the fundamental phonetic units that compose spoken words. Incorporating the context of phonemes into modeling offers a nuanced source of information for refining word embeddings. In our approach, we introduce a dedicated phoneme classifier into the training framework. Specifically, we adopt the AAM loss for phoneme classification. The phoneme loss is computed by aggregating the frame-level AAM loss on the phoneme labels across all frames.
2.4.4 Multi-Task Learning
@@
Consequently, the hybrid loss function for multi-task learning is constructed as a combination of word-level loss, the reverse speaker loss, and phoneme-level loss.
$\small\begin{split}\mathcal{L}(\mathbf{x},\mathbf{y})=\mathcal{L}\left(\mathbf{x},y^{w}\right)-\eta\mathcal{L}_{aam}(\mathbf{x},y^{s})+\mu\mathcal{L}_{aam}(\mathbf{x},y^{p}),\end{split}$
@@
where $\mathbf{x}$ is the acoustic feature vector, and $\mathbf{y}=(y^{w},y^{s},y^{p})$. $y^{w}$ is the word label, $y^{s}$ is the speaker label, and $y^{p}\in\mathbb{R}^{T}$ is the phoneme label sequence of $T$ frames. $\eta$ and $\mu$ are scaling factors. Note that the word-level loss $\mathcal{L}\left(\mathbf{x},y^{w}\right)$ can be expressed as either $\mathcal{L}_{aam}$ using AAM or $\mathcal{L}_{st}$ using SoftTriplet.
@@ 3.1 Dataset
We use the Librispeech [27] dataset containing 960 hours of read English audiobooks sampled at $16$ kHz along with transcriptions. We employ a pre-trained acoustic model for the force-alignment to segment utterances into individual words. Each word-level segment is standardized to $2$s long by clipping or zero padding on both sides of the audio. We use the internal aggregated and de-identified keyword dataset for evaluation. The positive data contains $275.7$k utterances from $629$ speakers. The total duration of negative data is up to $200$ hours. We extract acoustic features using $40$-dimensional log Mel-filterbank energies computed over a $25$ms window every $10$ms. The evaluation dataset employed in this study is entirely separate from the training data, ensuring that the keywords used for evaluation are not revealed in the training set. This setup aligns with the principles of customized keyword spotting. Additionally, the substantial size of the dataset ensures robust experimental results and conclusive findings.
@@ 3.2 Experimental Setup
Model architecture We conduct the experiments on three architectures as described in Section 2.2: LicoNet, Conformer, ECAPA_TDNN. We construct LiCoNet by stacking five LiCo-Blocks with the expansion factor of $6$ and the kernel size of $5$ [16]. Conformer has two heads per multi-headed self-attention layer with $128$ input and output nodes [28]. The linear hidden units have a dimensionality of $192$, and the convolution module uses the kernel size of $7$. We setup ECAPA_TDNN with $128$ channels in the convolution layers and a $64$ dimensional bottleneck in the SE-Block and attention module. The scale dimension in the Res2Block is $8$. Table 2 presents the model size and floating point operations per second (FLOPs) of $2$s audio for each encoder model.
@@
Feature aggregator We compare GAP against ASP as the feature aggregator. The spectral, temporal and spectro-temporal attention blocks use pooling ratios of $0.71$, $0.86$, and $0.71$.
Loss function We focus on investigating the effectiveness of different loss formulations. Due to limited space, we only show the best results from SoftTriplet. Similar improvements can also be seen in other setups. $m$ and s in Eq. 1 are set to $0.2$ and $32$. SoftTriplet uses $\lambda=60$, $\delta=0.03$, and $K=10$. The weights $\eta$ and $\mu$ in multi-task learning are set to $0.1$ and $0.5$.
@@
Training and testing protocols All KWS models are trained to predict 1002 targets (i.e., the top $1$k frequent words, Silence, and Unknown). We use a batch size of 64 with 8 GPUs for 40-epoch training. We adopt the triangular2 policy [29] using the Adam optimizer with a cyclical learning rate increased from $1$e-$8$ to $1$e-$3$ in $20$k warming-up updates. During testing, $3$ utterances of any speakers were randomly picked as enrollments. Given a query, the cosine distance is used to compare the similarity between embeddings of the query and the $3$ enrollments. The minimum distance is used to compare against a threshold to make the detection decision. We present the model performance by plotting detection error trade-off (DET) curves, where the x-axis and y-axis represent the number of false accepts (FA) per hour and false reject rate (FRR).
3.3 Results and Discussion
Table 1 summarizes FRR for each experiment at $0.3$ FAs/Hr. In Figure 2, we present DET curves for Conformer using various loss formulations and pooling strategies (left), and those for different encoders in the optimal multi-task learning setup (right). It is evident that all model architectures achieve their best performance in the hybrid loss configuration using word SoftTriplet and GAP. Particularly, LiCoNet demonstrates comparable performance to Conformer in the optimal setup.